Find the area of the surface obtained by rotating the circle x^2 +y^2 =3.2^2 about the line y = 3.2. (Keep three decimal places.)
Answer #1

som 6. ret Shift th У 23.? circle x²x₂33-2 В. це ъч 3.2 от +s and Rotate it about saadis will give the same result Now circlePut y 32= 3.2 sint dy=3.2 cost at Now y=3.2 +3-2 sint for uzo, t= -17 9 = 6:4, t = 2 Na => 5=1273 (3:2+3.2 sin t) x 3.2 cost